OpenStudy (anonymous):

does anyone know how 2n+2! becomes (2n+2)(2n+1)

OpenStudy (anonymous):

i guess it must b (2n +2)! ?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

n! = n(n-1)!

OpenStudy (anonymous):

(2n+2)(2n+1) (2n)....1

OpenStudy (anonymous):

(2n+2)!= (2n+2)(2n+1)!

OpenStudy (anonymous):

on cramster one of the problem shows it changes from (2n+2)! and becomes (2n+1)(2n+2) with no factorial sign behind it

OpenStudy (anonymous):

there is a facatorial behind it......

OpenStudy (anonymous):

i mean they don't show it like this (2n+2)(2n+1) (2n)....1

OpenStudy (anonymous):

I meant that u keep decreasing by 1 until u reach 1 another way of writing it (2n+2)(2n+1) (2n)...3.2.1

OpenStudy (anonymous):

yah they dont show it decreasing it stops after 2n+1

OpenStudy (anonymous):

yes but there must be a factorial with it

OpenStudy (anonymous):

and there isnt

OpenStudy (anonymous):

As far as I know about factorials ...there should be...I could be wrong

OpenStudy (anonymous):

is there anyway you can look at the problem on cramster

OpenStudy (anonymous):

u can check on wolframalpha.com

OpenStudy (anonymous):

there must be factorial with it

OpenStudy (anonymous):

there has to be a factorial behind it, unless there is something else about your problem... cramster sometimes has mistakes

OpenStudy (anonymous):

the things is my teacher did the same thing so im confused

OpenStudy (anonymous):

the only way the factorial can drop out is if it cancelled out in a fraction

OpenStudy (anonymous):

then check did u post the right question?

OpenStudy (anonymous):

this is the actual problem determine of this series converges or diverges\[\sum_{n=1}^{\infty} (n!)^2/(2n) !\]

OpenStudy (anonymous):

use ratio test

OpenStudy (anonymous):

n u applied some test for the convergence of series?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

i did but i didnt correct answer if someone can please show me how to do this it can help me with the rest of my problems

OpenStudy (anonymous):

sigh i hate it when people post half the question... thigns cancel and become (n+1)^2/(2n+2)(2n+1)

OpenStudy (anonymous):

what is 2n!

OpenStudy (anonymous):

yes it converges, so you use ratio test... an+1/an= [ ((n+1)!)^2/(2(n+1))! ]/ [ (n!)^2/(2n)! ] flipping fractions to get 1 whole one.. (2n)!(n+1)!(n+1)!/(n!)(n!)(2n+2)! which is where your question arises, like I said, the only reason the factorial disappeared is cause it got cancelled, which you always want to do when doing ratio test, so you simply keep pulling your factorials apart until they can cancel so on top (n+1)!=(n+1)(n!), which you do twice to get rid of both the n! on the bottom then on the bottom you have (2n+2)! so you pull the factorial apart twice to get (2n+2)(2n+1)(2n)! then the (2n)! can cancel with the one on top and you get left with (n+1)^2/(2n+2)(2n+1). take the limit and it should be 1/4 so it converges

OpenStudy (anonymous):

yes i was able to catch my mistake thank you very much. i wanted to know how does 2n! decrease

OpenStudy (anonymous):

does it become (2n)(2n-1)(2n-2)..1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey thank u for ur help im gonna become a fan